OpenStudy (kaylaprincess):

I am VERY confused to even know how to begin to solve this problem. I need someone to walk me through this problem. Medal as always. Two students in your class, Wilson and Alexis, are disputing a function. Wilson says that for the function, between x = -1 and x = 1, the average rate of change is 0. Alexis says that for the function, between x = -1 and x = 1, the graph goes up through a turning point, and then back down. Explain how Wilson and Alexis can both be correct, using complete sentences.

OpenStudy (kaylaprincess):

@ganeshie8 @Hero @Nnesha @iambatman much appreciation.

Nnesha (nnesha):

both can be wrong eh hun or nah ?

OpenStudy (kaylaprincess):

Well I'm not sure, it just says explain how both of them ~could~ be correct. For this should I use my knowledge of graphing functions and translation meanings or something different?

Nnesha (nnesha):

ohh nvm i thought both x are positive hm

Nnesha (nnesha):

well if slope is zero you will draw horizontal line

Nnesha (nnesha):

do you have graph with this question or table ?

OpenStudy (kaylaprincess):

No there is no graph or table.

Nnesha (nnesha):

i will use slope formula (average rate of change - slope )\[\huge\rm \frac{ y_2 - y_1}{ x_2 -x_1 }\] if slope is 0 then you should get 0 at the numerator substitute x's values x=1 and x=-1

OpenStudy (kaylaprincess):

What about Alexis?

OpenStudy (kaylaprincess):

Alexis says that for the function, between x = -1 and x = 1, the graph goes up through a turning point, and then back down. goes up through a turning point, then back down? ?

Nnesha (nnesha):

well how would you get 0 at the numerator in order to get zero both y values should be the same like y_1 = 2 y_2 = 2 then you will get 0 \[\frac{ 2-2 }{ 1-(-1) } = \frac{ 0 }{ 2}\] when slope is zero you will draw horizontal line  x values are increasing but y's is constant